# Asteria: Semantic-Aware Cross-Region Caching for Agentic LLM Tool Access

### v1

@@ Abstract
To address this challenge, we introduce Asteria, a novel cross-region knowledge caching architecture for LLM agents. At its core are two abstractions: Semantic Element (SE) and Semantic Retrieval Index (Sine). A semantic element captures the semantic embedding representation of an LLM query together with performance-aware metadata such as latency, cost, and staticity. Sine then provides two-stage retrieval: a vector similar index with semantic embedding for fast candidate selection and a lightweight LLM-powered semantic judger for precise validation. Atop these primitives, Asteria builds a new cache interface that includes a new semantic-aware cache hit definition, a cost-efficient eviction policy, and proactive prefetching. To reduce overhead, Asteria co-locates the small LLM judger with the main LLM using adaptive scheduling and resource sharing. Our evaluation demonstrates that Asteria delivers substantial performance improvements without compromising correctness. On representative search workloads, Asteria achieves up to a 3.6 increase in throughput by maintaining cache hit rates of over 85%, while preserving accuracy virtually identical to non-cached baselines. Asteria also improves throughput for complex coding tasks by 20%, showcasing its versatility across diverse agentic workloads.
@@ 1 Introduction
LLM agents [28, 66, 20] have emerged as powerful tools, capable of autonomously executing data-intensive tasks such as deep research assistance [33, 58, 39] and sophisticated code generation [14, 5, 44]. Their effectiveness relies on a multi-step reasoning and retrieval loop in which the agents strategically query external knowledge from either private knowledge bases via retrieval-augmented generation (RAG) [37] or cloud search services through API calls [26, 53], and then synthesize the retrieved information step by step to produce a final solution. Yet, this fundamental dependence on external interaction introduces a critical data-retrieval bottleneck to agent performance: the high latency and cost of accessing external knowledge for robust reasoning and decision-making.
@@
Compared to non-agent LLM inference, which relies mainly on GPU computation to generate text, LLM agents frequently invoke external data retrieval services/tools such as web search APIs [43, 57] or RAG backend services [30, 10]. Often, the data source and agent models are located in different data centers or regions and are connected by a wide-are network (WAN). Agent deployment thus faces two major challenges: high monetary cost and cross-region latency due to remote tool calling. For an application handling 5-10 million daily queries (such as Google AI mode [52, 27]), costs are substantial. Given the price of each Google Search API call at $0.005 [23], the application would incur a monthly costs of $1.5-4.5 million. Furthermore, accessing data sources in remote regions can lead to 300-500 ms end-to-end latencies [55]. Such high latency can impact user experience and disrupt multi-step agent workflows.
@@
Prior work has considered applying caching to reduce agent cost. However, existing caching strategies are designed to accelerate LLM inference itself, rather than mitigating the cost and latency of external data retrieval. Semantic prompt caches [9, 24, 65, 56] like GPT-Cache [9] store LLM outputs and reuse responses via prompt similarity to skip LLM generation. However, they compare prompts in the embedding space rather than at the knowledge or tool boundary, and they lack validation to ensure that reused content remains correct for the current context or time. Traditional data storage caches, such as key-value store [59], database [13] or file system [21], store KV objects indexed by exact keys. These caches lack the ability to evaluate semantic equivalence between non-matching queries. Transformer KV-caches [36, 22, 41] store token KV states to accelerate model decoding, but their scope is limited to inference computation.
@@
To address this gap, we propose a new paradigm, semantic-aware remote knowledge caching, to address the latency and cost bottlenecks in cross-region data access for LLM agents. Unlike prior solutions that focus on optimizing LLM inference computation (e.g., GPTCache) or relying on exact-match queries, our approach leverages LLMs’ language understanding to intelligently cache and retrieve external data. Knowledge caching is built on two core abstractions. We first define a semantic element (SE) that encapsulates the agent’s query, tool interactions, and the retrieved response. SE is augmented with performance-aware metadata such as latency, cost, and staticity. We then propose Semantic Retrieval Index (Sine), a two-stage retrieval engine that combines 1) an Approximate Nearest Neighbor (ANN) search for high-recall candidate selection and 2) a lightweight LLM-based Semantic Judger for precise validation, ensuring true semantic equivalence.
@@
We present Asteria, a concrete implementation of semantic-aware knowledge caching. Asteria seamlessly integrates the power of LLM semantic understanding into a robust caching architecture, significantly reducing external dependencies and costs associated with remote data access. Crucially, Asteria bridges the gap between the uncertainty of semantic matching and the deterministic requirements of a traditional cache. This is achieved by transforming the ANN and semantic judger output into a reliable, semantic-aware cache hit signal. This procedure ensures that only genuinely relevant and contextually appropriate remote information is served. Furthermore, Asteria equips the semantic index with advanced caching policies, such as a cost-efficient adaptive eviction policy and predictive prefetching, driven by SE metadata. Meanwhile, we propose an efficient Asteria implementation using GPU co-location, where the primary agent LLM and the semantic judger LLM can reside on a single GPU managed by a priority-aware scheduler that protects the agent’s critical latency paths.
@@
Our evaluation demonstrates that Asteria delivers substantial performance gains without compromising accuracy. On representative Zipfian and bursty workloads, Asteria achieves up to a 3.6 throughput improvement over exact-match caching by sustaining cache hit rates of over 85%. Crucially, our accuracy analysis reveals that while a naive semantic cache suffers a significant drop in correctness, Asteria maintains accuracy virtually identical to the non-cached baseline, proving the indispensable role of the semantic judger. This efficiency extends even to complex coding tasks, where Asteria provides a 20% throughput boost.
@@ 2.1 Agentic Applications Atop LLM
With the rise of reinforcement learning [17, 18], LLMs are empowered with thinking capability, moving beyond simple text generation to generate tool-use commands/queries in intermediate reasoning steps. The most mature ones are searching agent [32] and coding agents [5]
Agentic LLM. Unlike the traditional base LLM models [2], whose performance is primarily constrained by GPU inference speed, the core of an agentic system is an iterative “think-act-observe” loop that repeatedly accesses remote knowledge. As illustrated in 1(a), this process begins with a reasoning phase (①, think), where the agent formulates a plan. This plan then materializes when the LLM generates an action (i.e., tool call query) to be dispatched over the model context protocol (MCP [4]) to a remote data source. The tool call can be a search query or an API request. The dedicated think and action type of LLM outputs are encapsulated by special tags like <think> and <tool>. After retrieving the remote knowledge, the agent integrates it into its working context to inform the next cycle. This loop of external data retrieval is not a one-shot step but a fundamental, repeated operation at the heart of the agent’s problem-solving process, distinguishing its operational profile from non-agentic LLMs.
The Search-R1 [33] example. To make this workflow concrete, consider the behavior of a search-focused agent as depicted in 1(b). When tasked with finding the painter of the Mona Lisa, the agent first articulates its internal goal: <think> I need to find out who painted the Mona Lisa. </think>. It then translates this thought into a concrete action, issuing a search tool call with the query who painted the Mona Lisa?. This action triggers a remote tool call to an external search engine. Upon receiving the result, the agent observes it in an <info> block. Each search and info pair represents a single, costly round trip to an external data source.
@@
Heavy dependence on remote data sources creates significant latency and cost bottlenecks. Consider a standard cross-region deployment: an agent powered by a 7B model [33, 34] runs on a single H100 GPU in one region while its external tools (e.g., a Search API) reside in another, introducing 100–300 ms of network delay for every call.
Latency. External retrieval often rivals the model’s own inference time. As shown in 1(c), for running a search agent model Search-R1 workflow in a single H100, external data retrieval consistently constitutes around 40%-50% of the total execution time, forcing the expensive H100 GPU to remain idle for nearly half of the execution time and leading to a low GPU utilization of approximately 50%.
API rate limit. Beyond latency, agent serving throughput is also often throttled by external service limitations. For instance, commercial cloud APIs like Google Cloud Search [23] impose strict rate limits, i.e. 100 queries per minute per user. This creates a hard ceiling on request volume and introduces a severe bottleneck.
Financial cost. Furthermore, monetary impact can be equally severe. A service, such as Google AI mode, that handles 100 million MAU [52], with an average of 10 tool calls per query [51], would incur $0.005 per call, or roughly $150000 in daily API fees, as per Table 1. OpenAI charges $25 per 1000 tool calls for Web Search functionality [49], more expensively. For perspective, an H100 GPU costs about $1.49 per hour [1], so daily data-access charges alone can match the expense of 3300+ GPU-hours.
@@ 2.3 Access Pattern of the Agentic Workloads
Agentic LLM workloads rely heavily on remote tool calls, whose cost and latency often dominate performance. Caching can mitigate these overheads only if workloads exhibit sufficient locality. To assess this potential, we analyze two representative domains: AI-assisted search and code generation, both of which show statistical structures favorable to caching.
Search-oriented agent workloads. Agentic search workloads are ultimately shaped by real-world human interests, which already drive global web search and are becoming more prominent as major engines adopt AI-powered modes [42, 12]. Google Search, for instance, now incorporates agentic features—multi-source synthesis, conversational follow-ups, and contextual memory—that resemble the multi-step reasoning of LLM agents. Since commercial query logs are proprietary, we use public Google Trends data as a forward-looking proxy for the query distribution faced by large-scale agentic search systems.
@@
Code-agent workloads. While search-driven agents reflect global information demand, agentic LLMs are also widely used for code generation and software maintenance, which show a similarly skewed access pattern. To examine this, we analyzed coding tasks from SWE-Bench [47] on the sqlfluff [11] repository and measured how often each file is needed across tasks. As shown in Table 2, file accesses follow a near-Zipfian distribution: one file is required by nearly all tasks, a few core modules are reused heavily, and most files are rarely touched. This long-tail pattern parallels the popularity skew seen in search workloads. This also implies that caching these hot files locally can eliminate many redundant cross-region fetches, and semantic matching can further capture requests that refer to the same file.
@@ 2.4 Motivation and Cache Limitations
Motivation. Search-style and coding tasks show zipfian popularity with bursty reuse, suggesting that caching could greatly reduce redundant remote calls. Yet user queries in agentic workloads are semantically aligned rather than exactly key-matched, and they run in dynamic environments with practically complex factors like cross-region latency, retrieval cost, and tool API limits. A useful cache must handle semantic matching among queries while accounting for cost, cross-region latency, and tool rate limits.
@@
Existing caches and limitations. There are three common flavors of caching approaches, as summarized in Table 3, namely, transformer key-value (KV) caches [36, 22, 41], semantic prompt caches [9, 24, 65, 56], and traditional storage caches [59, 13] or file system [21]. Despite their popularity, the existing caching mechanisms exhibit significant drawbacks when coming to cache agentic workloads.
∗Semantic prompt cache performs vector similarity matching but lacks correctness validation.
@@
The first, transformer KV caches, accelerate token generation by reusing intermediate tensors for exactly-matched prompts, but its scope is confined to inference computation and does not extend to the remote data access layer. The second, semantic prompt cache, attempts to bypass inference entirely by matching user inputs via vector similarity. However, this approach suffers from an unfavorable precision-recall trade-off: vector proximity often fails to capture true semantic equivalence, so a query like “apple nutrition facts” might incorrectly match “Apple stock price analysis”. As a result, it is unreliable for production use and does not support semantic match completely. Also they only consider initial user prompt query and don’t focus on costly cross-region knowledge/tool call. As for the data storage cache [50] in databases or file systems, although many studies have explored cross-region scenarios, their exact-match mechanisms are fundamentally inadequate for natural language queries in agent workloads. These systems treat semantically equivalent queries as distinct keys, so even minor wording changes trigger cache misses and unnecessary remote calls. Additionally, none of these caches fully addresses rate limits impact in agent scenario, so agent throughput can still be throttled even when cache hits.
@@
Takeaway. Agentic workloads require a cache that performs semantic match at the knowledge or tool boundary, validates correctness, is cost-aware, cross-region and rate-limit aware so that hits yield real end-to-end gains. In contrast to these existing families, Asteria satisfies all properties in Table 3, motivating its design in the next section.

### v2

# Cortex: Achieving Low-Latency, Cost-Efficient Remote Data Access For LLM via Semantic-Aware Knowledge Caching

@@ Abstract
To address this challenge, we introduce Cortex, a novel cross-region knowledge caching architecture for LLM agents. At its core are two abstractions: Semantic Element (SE) and Semantic Retrieval Index (Seri). A semantic element captures the semantic embedding representation of an LLM query together with performance-aware metadata such as latency, cost, and staticity. Seri then provides a two-stage retrieval pipeline: a vector similar index with semantic embedding for fast candidate selection and a lightweight LLM-powered semantic judge for precise validation. Atop these primitives, Cortex builds a new cache interface that includes a new semantic-aware cache hit definition, a cost-efficient eviction policy, and proactive prefetching. To reduce overhead, Cortex co-locates the smaller LLM judge with the main LLM using adaptive scheduling and resource sharing. Our evaluation demonstrates that Cortex delivers substantial performance improvements without compromising correctness. On representative search workloads, Cortex achieves up to a 3.6 increase in throughput by maintaining cache hit rates of over 85% while preserving accuracy virtually identical to non-cached baselines. Cortex also improves throughput for coding tasks by 20%, showcasing its versatility across diverse agentic workloads.
@@ 1 Introduction
LLM agents [hong2023metagpt, zheng2024large, fei2024multimodal] have emerged as powerful tools that autonomously execute tasks such as deep research assistance [jin2025search, song2025r1, li2025search] and sophisticated code generation [cursor-code, claude-code, ma2025sql]. They rely on multi-step reasoning loops that strategically query external knowledge from either private knowledge bases via retrieval-augmented generation (RAG) [lewis2020retrieval] or cloud search services through API calls [gunjan2012search, ray2025survey], and they then synthesize the retrieved information step by step to produce a final solution. Yet, this fundamental dependence on external data retrieval introduces a critical bottleneck to agent performance: high latency and cost of external knowledge access.
@@
Compared to non-agent LLM inference, which relies mainly on GPU computation to generate text, LLM agents frequently invoke external data retrieval services/tools such as web search APIs [googleSearch2025, serpapi2025] or RAG backend services [turbopuffer2025, chroma2025]. Often, the data source and agent models are located in different data centers or regions and are connected by a wide-area network (WAN). Agent deployment thus faces two major challenges: high monetary cost and cross-region latency due to remote tool calling. For applications [googleaimode-report, googleaimode-estimate] handling 5-10 million daily queries, costs are substantial. Given the price of each Google Search API call at $0.005 [searchapiprice], the application would incur a monthly costs of $1.5-4.5 million. Furthermore, accessing data sources in remote regions can lead to 300-500 ms end-to-end latencies [cross-region-latency]. Such high latency can impact user experience and disrupt multi-step workflows.
@@
Prior works have considered applying caching to reduce agent cost. However, existing caching strategies are designed to accelerate LLM inference itself, rather than mitigating the cost and latency of external data retrieval. Semantic prompt caches [bang2023gptcache, google-cache, yan2025contextcache, schroeder2025adaptive] like GPT-Cache [bang2023gptcache] store LLM outputs and reuse responses via prompt similarity to skip LLM generation. However, they compare prompts in the embedding space rather than at the knowledge or tool boundary, and they lack validation to ensure that reused content remains correct for the current context or time. Traditional data storage caches, such as key-value store [rocksdb2025], database [mysql2025] or file system [ceph2025], store KV objects indexed by exact keys. These caches lack the ability to evaluate semantic equivalence between non-matching queries. Transformer KV-caches [kwon2023efficient, gim2024prompt, liu2024cachegen] store token KV states to accelerate model decoding, but their scope is limited to inference computation.
@@
To address this gap, we propose a new paradigm, semantic-aware remote knowledge caching, to address latency and cost bottlenecks in cross-region data access for LLM agents. Unlike prior solutions that focus on optimizing LLM inference computation (e.g., GPTCache) or relying on exact-match queries, our approach leverages LLMs’ language understanding to intelligently cache and retrieve external data. Knowledge caching is built on two core abstractions. We first define a semantic element (SE) that encapsulates the agent’s query, tool interactions, and the retrieved response. SE is augmented with performance-aware metadata such as latency, cost, and staticity. We then propose Semantic Retrieval Index (Seri), a two-stage retrieval engine that combines 1) an Approximate Nearest Neighbor (ANN) search for high-recall candidate selection and 2) a lightweight LLM-based Semantic Judge for precise validation, ensuring true semantic equivalence.
@@
We present Cortex, a concrete implementation of semantic-aware knowledge caching. Cortex seamlessly integrates the power of LLM semantic understanding into a robust caching architecture, significantly reducing external dependencies and costs associated with remote data access. Crucially, Cortex bridges the gap between the uncertainty of semantic matching and the deterministic requirements of a traditional cache. This is achieved by transforming the ANN and semantic judge output into a reliable, semantic-aware cache hit signal. This procedure ensures that only genuinely relevant and contextually appropriate remote information is served. Furthermore, Cortex equips the semantic index with advanced caching policies, such as a cost-efficient adaptive eviction policy and predictive prefetching, driven by SE metadata. Meanwhile, we propose an efficient Cortex implementation using GPU co-location, where the primary agent LLM and the semantic judge LLM can reside on a single GPU managed by a priority-aware scheduler that protects the agent’s critical latency paths.
@@
Our evaluation demonstrates that Cortex delivers substantial performance gains without compromising accuracy. On representative Zipfian and bursty workloads, Cortex achieves up to a 3.6 throughput improvement over exact-match caching by sustaining cache hit rates of over 85%. Crucially, our accuracy analysis reveals that. while a naive semantic cache suffers a significant drop in correctness, Cortex maintains accuracy virtually identical to the non-cached baselines, proving the indispensable role of the semantic judge. This efficiency extends even to complex coding tasks, where Cortex provides a 20% increase in throughput.
@@ 2.1 Agentic Applications Atop LLM
With the rise of reinforcement learning [deepseekai2025deepseekv3technicalreport, kimiteam2025kimik2openagentic], LLMs are empowered with thinking capability, moving beyond simple text generation to generate tool-use commands/queries in intermediate reasoning steps. The most mature ones are searching agents [search_r1] and coding agents [claude-code].
Agentic LLM. Unlike base LLM models [qwen25model], primarily constrained by GPU inference speed, the core of an agentic system is an iterative “think-act-observe” loop that repeatedly accesses remote knowledge. As illustrated in 1(a), this process begins with a reasoning phase (①, think), where the agent formulates a plan. This plan then materializes when the LLM generates an action (i.e., tool call query) to be dispatched over the model context protocol (MCP [antropic2024mcp]) to a remote data source. The tool call can be a search query or an API request. The dedicated think and action type of LLM outputs are encapsulated by special tags like <think> and <tool>. After retrieving remote knowledge, the agent integrates it into its working context to inform the next cycle. This loop of external data retrieval is not a one-shot step but a repeated operation at the heart of the agent’s problem-solving process, distinguishing its operational profile from non-agentic LLMs.
The Search-R1 [jin2025search] example. To make this workflow concrete, consider the behavior of a search-focused agent as depicted in 1(b). When tasked with finding the painter of the Mona Lisa, the agent first articulates its internal goal: <think> I need to find out who painted the Mona Lisa. </think>. It then translates this thought into a concrete action, issuing a search tool call with the query who painted the Mona Lisa?. This action triggers a remote tool call to an external search engine. Upon receiving the result, the agent observes it in an <info> block. Each search and info pair represents a single, costly round trip to an external data source.
@@
| Company | OpenAI | ||
| Operation | Search API | Web Search Preview | Web Search |
| Cost (per 1k reqs.) | $5 [searchapiprice] | $10-$25 [openaiapipricing] | $10 [openaiapipricing] |
Heavy dependence on remote data sources creates significant latency and cost implications. Consider a standard cross-region deployment: an agent powered by a 7B model [jin2025search, jin2025model] runs on a single H100 GPU in one region while its external tools (e.g., a Search API) reside in another, introducing 100–300 ms network delay for every call.
Latency. External retrieval often rivals the model’s own inference time. As shown in 1(c), running Search-R1 on a single H100, external data retrieval constitutes around 40%-50% of the total execution time. While batching multiple requests can improve GPU utilization, the latency penalty of remote tool call is fundamental: each individual request must still wait for its remote tool calls to complete, as these calls lie on the critical path of the agent’s reasoning loop. This creates a latency floor that cannot be reduced via parallelism alone, directly impacting user-perceived response time.
Financial cost. For large-scale deployments, API fees can become a substantial operational expense. Take Google AI mode that handles 100 million MAU [googleaimode-report] as an example. With an average of 10 tool calls per query [patil2023gorilla], the service would incur $0.005 per call, or roughly $150,000 in daily API fees, as per Table 1. OpenAI charges up to $25 per 1000 tool calls for Web Search functionality [openaiapipricing], which is more expensive. To serve such query volume with a production-grade dense model (e.g., Qwen2.5-72B or Llama-3-70B), a deployment requires approximately 50–100 H100 GPUs (at roughly $1.49/hour each [hyperbolic2025], or $1,800–$3,600/day). For this service, the $150,000 daily API cost is 40–80 higher than that of GPU compute, making external data access the dominant operational expense.
API rate limit. Beyond latency and cost, external API rate limits raise practical challenges for smaller-scale deployments. For instance, commercial cloud APIs like Google Cloud Search [searchapiprice] impose strict rate limits, e.g., 100 queries per minute per user. Large enterprises can circumvent this by maintaining multiple accounts or negotiating higher quotas, but such solutions are impractical for individual developers and small teams. As a result, many resort to third-party aggregation services (e.g., SerpAPI), which add both cost and external dependencies. Cortex offers a lightweight alternative that reduces the number of external calls in the first place.
@@ 2.3 Access Pattern of the Agentic Workloads
Agentic LLM workloads heavily rely on remote tool calls, whose cost and latency often dominate performance. Caching can mitigate these overheads only if workloads exhibit sufficient locality. To assess this potential, we analyze two representative domains: AI-assisted search and code generation, both of which show statistical structures favorable to caching.
Search-oriented agent workloads. Agentic search workloads are ultimately shaped by real-world human interests, which already drive global web search and are becoming more prominent as major engines adopt AI-powered modes [googleaimode2025, newbing2025]. Google Search, for instance, now incorporates agentic features—multi-source synthesis, conversational follow-ups, and contextual memory—that resemble the multi-step reasoning of LLM agents. Since commercial query logs are proprietary, we use public Google Trends data as a forward-looking proxy for the query distribution faced by large-scale agentic search systems.
@@
Code-agent workloads. While search-driven agents reflect global information demand, agentic LLMs are also widely used for code generation and software maintenance, which show a similarly skewed access pattern. To examine this, we analyzed coding tasks from SWE-Bench [swebench_oracle] on the sqlfluff [sqlfluff_repo] repository and measured how often each file is needed across tasks. As shown in Table 2, file accesses follow a near-Zipfian distribution: one file is required by nearly all tasks, a few core modules are reused heavily, and most files are rarely touched. This long-tail pattern parallels the popularity skew seen in search workloads. This also implies that caching these hot files locally can eliminate many redundant cross-region fetches, and semantic matching can further capture requests that refer to the same file.
@@ 2.4 Motivation and Cache Limitations
Motivation. Search-style and coding tasks show Zipfian popularity with bursty reuse, suggesting that caching could greatly reduce redundant remote calls. Yet, user queries in agentic workloads are semantically aligned rather than exactly key-matched, and they run in dynamic environments with complexities like cross-region latency, retrieval cost, and tool API limits. A useful cache must handle semantic matching among queries while accounting for cost, cross-region latency, and tool rate limits.
@@
Existing caches and limitations. There are three common flavors of caching approaches, as summarized in Table 3, namely, transformer key-value (KV) caches [kwon2023efficient, gim2024prompt, liu2024cachegen], semantic prompt caches [bang2023gptcache, google-cache, yan2025contextcache, schroeder2025adaptive], and caches for traditional storage [rocksdb2025, mysql2025] or file system [ceph2025]. Despite their popularity, existing caching mechanisms exhibit significant drawbacks when handling agentic workloads.
| System | Cached data | Semantic Match | Cost-aware Eviction | Cross-region Aware | Rate-limit aware |
| Transformer KV-cache [kwon2023efficient, gim2024prompt, liu2024cachegen] | Token KV states | ✗ | ✗ | ✗ | ✗ |
| Semantic prompt cache [bang2023gptcache, google-cache, yan2025contextcache, schroeder2025adaptive] | LLM response | ✗ | ✗ | ✗ | ✗ |
| Traditional storage cache [rocksdb2025, mysql2025, ceph2025] | KV object | ✗ | ✓ | ✓ | ✗ |
| Cortex | External knowledge | ✓ | ✓ | ✓ | ✓ |
@@
The first type, transformer KV caches, accelerates token generation by reusing intermediate tensors for exactly-matched prompts, but its scope is confined to inference computation and does not extend to the remote data access layer. The second type, semantic prompt caches (e.g., ContextCache [yan2025contextcache], VectorQ [schroeder2025adaptive]), attempts to bypass inference entirely by matching user inputs via vector similarity. However, this approach suffers from a precision-recall trade-off: vector proximity often fails to capture true semantic equivalence. For example, VectorQ, despite applying adaptive thresholds for similar prompts, still relies on the limited semantic awareness of vector similarity, so a query like “apple nutrition facts” might incorrectly match “Apple stock price analysis”. Moreover, these approaches are designed for plain LLM inference rather than agentic workloads. They cache model outputs at the prompt level but do not intercept or optimize external tool calls. For instance, ContextCache targets conversational LLMs and is not designed to handle the tool-calling patterns. Lastly, studies have explored the data storage cache [park2024reducing] in databases or file systems in cross-region use cases. However, their exact-match mechanisms are fundamentally inadequate for natural language queries in agent workloads. These systems treat semantically equivalent queries as distinct keys, so even minor wording changes trigger cache misses and unnecessary remote calls. Additionally, none of these caches fully addresses the rate limit impact in the agent scenario, so agent throughput can still be throttled even when cache hits.
@@
Takeaway. Agentic workloads require a cache that performs semantic match at the knowledge or tool boundary, validates correctness, is cost-aware, cross-region, and rate-limit aware so that cache hits yield real end-to-end gains. In contrast to these existing families, Cortex satisfies all required properties as shown in Table 3, motivating its design in the next section.